Given a demand curve Qd = 4000 – 250P and supply curve Qs = -200 + 500P, what are the equilibrium price and quantity levels that could be expected for this product? (15%) How would the demand function be affected by each of the following actions? (15%) An increase in price by $1.00, of the product above the equilibrium price as calculated above. A decrease in the number of suppliers of the product? A reduction in personal income of buyers in the market? Strong evidence of higher demand among millennials because of a recent advertising campaign? The price of a competitive product announced a 30% increase in prices? Considering the equilibrium price/quantity calculations and associated adjustments as outlined above, please provide a brief narrative of the potential success of this product? (10%)
Answer #1

Qd(Quantity demanded)= 4000-250P and Quantity supplied Qs = -200+500P

Equilibrium is when Qd= Qs

4000-250P= -200+500P

4200=750P

P= 5.6 and Quantity is : 4000-250*5.6 = 2600

How demand function is affected:

when:

An increase in price by $1.00, of the product above the equilibrium price as calculated above.

New price is $6.6 and hence demand quantity will be: 4000-250(*6.6)= 2350 units, demand decreases by : 250 units.

A decrease in the number of suppliers of the product?

Decrease in suppliers will make price to go up and as supply curve shifts left.

A reduction in personal income of buyers in the market?

This will shift demand curve to left and prices will come down.

Strong evidence of higher demand among millennials because of a recent advertising campaign? The price of a competitive product announced a 30% increase in prices? Considering the equilibrium price/quantity calculations and associated adjustments as outlined above, please provide a brief narrative of the potential success of this product? (10%)

It can be expected that demand is inelastic as price increase by 17.85 5 led to decrease in quantity only by 9.61 % this shows inelastic demand.

Competitor increasing prices will not decrease its revenue. This product has to rely on its own markets, advertisement campaigns to increase its profitability.